OpenStudy (anonymous):

What is the quivalent of √3 +5 /√3 -5

OpenStudy (anonymous):

is it \[\frac{\sqrt{3}+5}{\sqrt{3}-5}\]?

OpenStudy (anonymous):

no,i think you have to multiply the whole thing by √3 +5 but I don't remember how

OpenStudy (anonymous):

no i meant was that the question?

OpenStudy (anonymous):

oh yeah, sorry

OpenStudy (anonymous):

if so multiply numerator and denominator by the conjugate of the denominator

OpenStudy (anonymous):

\[\frac{\sqrt{3}+5}{\sqrt{3}-5}\times \frac{\sqrt{3}+5}{\sqrt{3}+5}\]

OpenStudy (anonymous):

the denominator will be \[3-25=-22 \]

OpenStudy (anonymous):

the numerator will be \[3+25+10\sqrt{3}\]

OpenStudy (anonymous):

is the final, simplified answer 14 + 5√3 / -11 ?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

ok thank you!

OpenStudy (anonymous):

welcome!